\begin{thebibliography}{10}

\bibitem{AvgSol:1996:nonsyst}
S.~V. Avgustinovich, F.~I. Solov'eva.
\newblock On the nonsystematic perfect binary codes.
\newblock {\em \href{http://link.springer.com/journal/11122}{Probl. Inf.
  Transm.}}, 32(3):258--261, 1996.
  Translated from \em
  \href{http://www.mathnet.ru/php/journal.phtml?jrnid=ppi\&option_lang=eng}{Probl.
  Peredachi Inf.}\em , 32(3):47--50, 1996.

\bibitem{AvgSol:1998IT:trivAut}
S.~V. Avgustinovich, F.~I. Solov'eva.
\newblock Perfect binary codes with trivial automorphism group.
\newblock In {\em Proc. the {I}nformation Theory Workshop}, pages 114--115,
  Killarney, Ireland, June 1998.
\newblock \DOI{10.1109/ITW.1998.706460}.

\bibitem{ASH:RankKernel}
S.~V. Avgustinovich, F.~I. Solov'eva, O.~Heden.
\newblock On ranks and kernels problem of perfect codes.
\newblock {\em \href{http://link.springer.com/journal/11122}{Probl. Inf.
  Transm.}}, 39(4):341--345, 2003.
\newblock \DOI{10.1023/B:PRIT.0000011272.10614.8c} translated from \em
  \href{http://www.mathnet.ru/php/journal.phtml?jrnid=ppi\&option_lang=eng}{Probl.
  Peredachi Inf.}\em , 39(4):30--34, 2003.

\bibitem{BryHor:Lindner}
D.~Bryant, D.~Horsley.
\newblock A proof of {L}indner's conjecture on embeddings of partial {S}teiner
  triple systems.
\newblock {\em
  \href{http://onlinelibrary.wiley.com/journal/10.1002/(ISSN)1520-6610}{J.
  Comb. Des.}}, 17(1):63--89, 2008.
\newblock \DOI{10.1002/jcd.20189}.

\bibitem{ColRos}
C.~J. Colbourn, A.~Rosa.
\newblock {\em Triple Systems}.
\newblock Clarendon Press, Oxford, 1999.

\bibitem{EV:94}
T.~Etzion, A.~Vardy.
\newblock Perfect binary codes: Constructions, properties and enumeration.
\newblock {\em
  \href{http://ieeexplore.ieee.org/xpl/RecentIssue.jsp?punumber=18}{IEEE Trans.
  Inf. Theory}}, 40(3):754--763, 1994.
\newblock \DOI{10.1109/18.335887}.

\bibitem{Ganter:QStoSQS}
B.~Ganter.
\newblock Finite partial quadruple systems can be finitely embedded.
\newblock {\em
  \href{http://www.sciencedirect.com/science/journal/0012365X}{Discrete
  Math.}}, 10(2):397--400, 1974.
\newblock \DOI{10.1016/0012-365X(74)90130-7}.

\bibitem{Hed:2008:survey}
O.~Heden.
\newblock A survey of perfect codes.
\newblock {\em \href{http://aimsciences.org/journals/amc/}{Adv. Math.
  Commun.}}, 2(2):223--247, 2008.
\newblock \DOI{10.3934/amc.2008.2.223}.

\bibitem{KroAvg:2008:lb}
D.~S. Krotov, S.~V. Avgustinovich.
\newblock On the number of $1$-per\-fect binary codes: A lower bound.
\newblock {\em
  \href{http://ieeexplore.ieee.org/xpl/RecentIssue.jsp?punumber=18}{IEEE Trans.
  Inf. Theory}}, 54(4):1760--1765, 2008.
\newblock \DOI{10.1109/TIT.2008.917692}.

\bibitem{Linder:survey-embed}
C.~C. Lindner.
\newblock A survey of embedding theorems for {S}teiner systems.
\newblock In C.~C. Lindner and A.~Rosa, editors, {\em Topics on Steiner
  Systems}, volume~7 of {\em Ann. Discrete Math.}, pages 175--202.
  North-Holland, 1980.

\bibitem{Mal:1998OC:trivAut}
S.~A. Malyugin.
\newblock Perfect codes with trivial automorphism group.
\newblock In {\em Proc. the {I}nt. Workshop on Optimal Codes}, pages 163--167,
  Sozopol, Bulgaria, June 1998.

\bibitem{Mal:2001:nonsyst}
S.~A. Malyugin.
\newblock Nonsystematic perfect binary codes.
\newblock {\em
  \href{http://www.mathnet.ru/php/journal.phtml?jrnid=da\&option_lang=eng}{Diskretn.
  Anal. Issled. Oper.}, Ser.~1}, 8(1):55--76, 2001.
\newblock In Russian.

\bibitem{OstPot2007}
P.~R.~J. {\"O}sterg{\aa}rd, O.~Pottonen.
\newblock There exist {S}teiner triple systems of order $15$ that do not occur
  in a perfect binary one-error-correcting code.
\newblock {\em
  \href{http://onlinelibrary.wiley.com/journal/10.1002/(ISSN)1520-6610}{J.
  Comb. Des.}}, 15(6):465--468, 2007.
\newblock \DOI{10.1002/jcd.20122}.

\bibitem{PheLeV:1995}
K.~T. Phelps, M.~LeVan.
\newblock Kernels of nonlinear {H}amming codes.
\newblock {\em \href{http://link.springer.com/journal/10623}{Des. Codes
  Cryptography}}, 6(3):247--257, 1995.
\newblock \DOI{10.1007/BF01388478}.

\bibitem{PheLeV:1999}
K.~T. Phelps, M.~LeVan.
\newblock Nonsystematic perfect codes.
\newblock {\em \href{http://epubs.siam.org/journal/sjdmec}{SIAM J. Discrete
  Math.}}, 12(1):27--34, 1999.
\newblock \DOI{10.1137/S0895480196312206}.

\bibitem{PheVil:RankKernel}
K.~T. Phelps, M.~Villanueva.
\newblock On perfect codes: Rank and kernel.
\newblock {\em \href{http://link.springer.com/journal/10623}{Des. Codes
  Cryptography}}, 27(3):183--194, 2002.
\newblock \DOI{10.1023/A:1019936019517}.

\bibitem{Sol:2008:survey}
F.~I. Solov'eva.
\newblock On perfect binary codes.
\newblock {\em
  \href{http://www.sciencedirect.com/science/journal/0166218X}{Discrete Appl.
  Math.}}, 156(9):1488--1498, 2008.
\newblock \DOI{10.1016/j.dam.2005.10.023}.

\bibitem{Treash:TStoSTS}
C.~Treash.
\newblock The completion of finite incomplete {S}teiner triple systems with
  applications to loop theory.
\newblock {\em \href{http://www.sciencedirect.com/science/journal/00973165}{J.
  Comb. Theory, Ser.~A}}, 10(3):259--265, 1971.
\newblock \DOI{10.1016/0097-3165(71)90030-6}.

\bibitem{Vas:nongroup_perfect}
Yu.~L. Vasil'ev.
\newblock On nongroup close-packed codes.
\newblock In {\em Probleme der Kybernetik}, volume~8, pages 92--95. 1965.
\newblock Translated from {\em Problemy Kibernetiki}, volume~8, pages 337--339. 1962.

\bibitem{VAK:2008}
Yu.~L. Vasil'ev, S.~V. Avgustinovich, D.~S. Krotov.
\newblock On shifting sets in the binary hypercube.
\newblock \href{http://link.springer.com/journal/11754}{J. Appl. Ind. Math.},
3(2):290--296, 2009.
Translated from
\newblock {\em
  \href{http://www.mathnet.ru/php/journal.phtml?jrnid=da\&option_lang=eng}{Diskretn.
  Anal. Issled. Oper.}}, 15(3):11--21, 2008.
\newblock arXiv:\href{http://arxiv.org/abs/0802.0003}{0802.0003}

\end{thebibliography}

\documentclass[12pt,letterpaper]{article}
\usepackage[english]{babel}
\usepackage{hyperref}

\date{}
\newcommand\proofr[1][\hspace{-1em}]{\par{\em Proof\hspace{1em}#1:~}}

\def\proofend{$\bigtriangleup$\vspace{0.3em}\par}

\newcounter{theorem}
\newenvironment{theorem}[1][\hspace{-1.0ex}]%
 {\par\addvspace{2mm}\indent\refstepcounter{theorem} \textbf{Theorem\hspace{1.0ex}{\rm#1}.~}\sl}%
 {\par\addvspace{2mm}\rm}
\newcounter{remark}
\newenvironment{remark}[1][\hspace{-1.0ex}]%
 {\par\addvspace{2mm}\indent\refstepcounter{remark} \textbf{Remark~\theremark\hspace{1.0ex}{\rm#1}.~}}%
 {\par\addvspace{2mm}\rm}
\newcounter{lemma}0
\newenvironment{lemma}[1][\hspace{-1.0ex}]%
 {\par\addvspace{2mm}\indent\refstepcounter{lemma} \textbf{Lemma~\thelemma\hspace{1.0ex}{\rm#1}.~}\sl}%
 {\par\addvspace{2mm}\rm}

\providecommand\href[2]{#2}
\author{Sergey V. Avgustinovich, Denis S. Krotov}

\def\zeroext#1{\bar #1}
\begin{document}


\begin{center}
{\LARGE Embedding in a perfect code%
\footnote{
This is the peer reviewed version of the following article: 
``Embedding in a perfect code'', Journal of Combinatorial Designs 17(5) 2009, 419--423, 
which has been published in final form at 
\url{http://dx.doi.org/10.1002/jcd.20207}.
This article may be used for non-commercial purposes in accordance with 
Wiley Terms and Conditions for Self-Archiving.
\\
The results of the paper were presented at the Workshop
``Coding Theory Days in St. Petersburg'', Oct. 2008, St. Petersburg, Russia.
\\
The authors are with the Sobolev Institute of Mathematics, Novosibirsk, Russia.
E-mail: \{avgust,krotov\}@math.nsc.ru
\\
The research was partially supported by the RFBR grant 07-01-00248 and 08-01-00673}
}
\end{center}

\begin{center}
{\large
{Sergey~V.~Avgustinovich},
{Denis~S.~Krotov}}
\end{center}

\renewcommand\abstractname{}
\begin{abstract}
\begin{center}
A binary $1$-error-correcting code can always be embedded \\in a $1$-perfect code
of some larger length.\end{center}
\end{abstract}

For any  $1$-error-correcting binary code $C$ of length $m$ we will construct a $1$-perfect
binary code $P(C)$ of length $n=2^m-1$ such that fixing the last $n-m$ coordinates by zeroes in $P(C)$
gives $C$.

In particular, any complete or partial Steiner triple system (or any other system
that forms a $1$-code)
can always be embedded in a $1$-perfect code of some length
(compare with \cite{OstPot2007}).
Since the weight-$3$ words of a $1$-perfect code $P$ with $0^n \in P$ form a Steiner triple system,
and the weight-$4$ words of an extended $1$-perfect code $P$ with $0^n \in P$
form a Steiner quadruple system, we have, as corollaries, the following well-known facts:
a patrial Steiner triple (quadruple) system can always be embedded
in a Steiner triple (quadruple) system \cite{Treash:TStoSTS} (\cite{Ganter:QStoSQS})
(these results, as well as many other embedding theorems for Steiner systems,
can be found in \cite{Linder:survey-embed,ColRos}).

Notation:
\begin{itemize}
\item $F^m$ denotes the set of binary $m$-tuples, or binary $m$-words.
\item $\dot F^m := F^m \setminus \{0^m \}$, where $0^m$ is the all-zeroes $m$-word.
\item$F^m$ is considered as a vector space over $GF(2)$ with calculations modulo $2$.
\item${{\Pi}}=\{\pi^{(1)},\ldots,\pi^{(m)}\}=\{(10..0),\ldots,(0..01)\}$ is the natural basis in $F^m$.
\item$n:=2^m-1$.
\item The elements of $F^m$ will be denoted by Greek letters.
\item The elements of $F^n$ will be denoted by overlined letters,
their coordinates being indexed by the elements of $\dot F^m$, e.g.,
$\bar w = \{w_\iota\}_{\iota\in \dot F^m}$; we assume
that the first $m$ coordinates have the indexes $\pi^{(1)},\ldots,\pi^{(m)}$, and
the order of the other $n-m$ indexes does not matter (but fixed).
\item
$\{\bar e^{(\iota)}\}_{\iota\in \dot F^m}$ is the natural basis in $F^n$;
note that $\bar e^{(\pi^{(\iota)})}=(\pi^{(\iota)},0^{n-m})$.
\item
For any $\alpha=(\alpha_1,...,\alpha_m)\in F^m$ denote
$$ \zeroext{\alpha} := (\alpha,0^{n-m});$$
it also holds $ \zeroext{\alpha} =\sum_{i=1}^{m} \alpha_i \bar e^{(\pi^{(i)})}. $
\item
$d(\cdot,\cdot)$ denotes the Hamming \emph{distance} between two words in $F^m$ or $F^n$
(the number of positions in which the words differ).
\item
$<\ldots>$ denotes the linear span of the vectors or sets of vectors between the angle brackets.
\item
The \emph{neighborhood} $\Omega(M)$ of a set $M\subset F^n$ is the set of vectors
at distance at most $1$ from $M$.
\item
A set $C\subset F^m$ is called a \emph{$1$-code} if the neighborhoods of the codewords are disjoint.
\item
A $1$-code $P\subset F^n$ is called a \emph{$1$-perfect code} if $\Omega(P)=F^n$;
in this case, $|P|=2^n/(n+1)$.
\item
The
\emph{Hamming code} $H$ defined as
\begin{equation}\label{eq:H}
H := \{ \bar c\in\{0,1\}^n | \sum_{\alpha\in\dot F^m} c_\alpha \alpha = 0^m   \}
\end{equation}
is a linear $1$-perfect code.
\item
For any $\iota$ from $\dot F^m$ the \emph{linear $\iota$-component} of $H$ is defined as
$$
R_\iota := \{ \bar c\in H |  c_\alpha = c_{\alpha+\iota} \mbox{ for all }  \alpha \in F^m\setminus <\iota>\}
$$
\end{itemize}
(note that $R_\iota$ is a linear subcode of $H$, for all $\iota$).
Since \cite{Vas:nongroup_perfect}, linear components are used for constructing
non-linear $1$-perfect codes.
For the first time, the method of synchronous switching nonintersecting
linear $i$-components with different $i$,
which is exploited in this paper
(to follow our notations, we replace $i$ by Greek letters),
was used in \cite{EV:94}.
This method was also used by different authors for constructing $1$-perfect codes with
specified properties,
such as different ranks and(or) kernels
\cite{EV:94,PheLeV:1995,PheVil:RankKernel,ASH:RankKernel},
trivial automorphism group \cite{AvgSol:1998IT:trivAut,Mal:1998OC:trivAut},
nonsystematicity \cite{AvgSol:1996:nonsyst, PheLeV:1999, Mal:2001:nonsyst}
(to read more about these and other results in the theory of $1$-perfect codes,
see \cite{Hed:2008:survey,Sol:2008:survey}).
The term ``switching'' was introduced in
\cite{PheLeV:1999}; probably, it came from the theory of Steiner triple systems.
In general, this term refers to replacing some part of a $1$-perfect code by some set with
the same cardinality and same neighborhood; the resulting code will also be $1$-perfect.
With this approach, a large number of $1$-perfect codes can be produced \cite{KroAvg:2008:lb}.
$1$-Codes that admit replacing by other $1$-codes with the same neighborhood
deserve the independent study; they exist in $F^n$ for every odd $n$ \cite{VAK:2008}.

%
Being such a $1$-code is the main property of $R_\iota$.
Since our definition of linear components differs from others,
we should prove this fact
(in essence, the following lemma coincides with \cite[Corollary 3.4]{EV:94}).

\begin{lemma}\label{l:O} For any $\bar z$ from $F^n$ it holds that
  $\Omega(R_\iota+\bar z) = \Omega(R_\iota+\bar z+\bar e^{(\iota)}).$
\end{lemma}
\proofr Without loss of generality, assume $\bar z = 0^n$.
Denote $\bar e^{(0^m)}:=0^n$. Then, we have
$$
\Omega(R_\iota)
=\bigcup_{\kappa\in F^m}(R_\iota + \bar e^{(\kappa)})
=\bigcup_{\kappa\in F^m}(R_\iota + \bar e^{(\iota)} + \bar e^{(\kappa+\iota)})
=\bigcup_{\lambda\in F^m}( (R_\iota + \bar e^{(\iota)}) + \bar e^{(\lambda)})
=\Omega(R_\iota + \bar e^{(\iota)})
$$
because $\bar e^{(\iota)} + \bar e^{(\kappa)} + \bar e^{(\kappa+\iota)} \in R_\iota$ for all $\kappa\in F^m$.
\proofend

\begin{lemma}\label{l:RR} Every element $\bar c$ of $<R_\iota,R_\kappa>$ satisfies
\begin{equation}\label{eq:RR}
 c_{\alpha} + c_{\alpha+\iota}+c_{\alpha+\kappa} + c_{\alpha+\iota+\kappa}=0
\mbox{ for all } \alpha \in F^m\setminus <\iota,\kappa>
\end{equation}
\end{lemma}
\proofr
By the definition, the elements of $R_\iota$ and $R_\kappa$ satisfy (\ref{eq:RR}).
Thus, the elements of their linear span also satisfy (\ref{eq:RR}).
\proofend

The following lemma is the crucial part of our reasoning.

\begin{lemma}\label{l:disj}
  For any $\iota,\kappa\in\dot F^m$ at distance at least $3$ from $0^m$
  and from each other,
  the $\iota$-component
  $R_\iota+ \zeroext{\iota} + \bar e^{(\iota)}$
  and
  the $\kappa$-component
  $R_\kappa+ \zeroext{\kappa} + \bar e^{(\kappa)}$
  are disjoint
  and do not contain $0^n$.
\end{lemma}
\proofr
By general algebraic reasons, it is enough to show that
$\bar w:= \zeroext{\iota} + \bar e^{(\iota)} + \zeroext{\kappa} + \bar e^{(\kappa)}$,
 does not belong to
$<R_\iota,R_\kappa>$.
Let $j$ be a nonzero coordinate of $\iota+\kappa$.
Then $\pi^{(j)}$ is the index of
a nonzero coordinate of $\bar w$;
the indexes of the other nonzero coordinates
also belong to ${\Pi} \cup \{\iota,\kappa\}$.
But, since the mutual distances between $0^m$, $\iota$, $\kappa$, and $\iota+\kappa$
are not less than $3$,
the indexes $\pi^{(j)}+\iota$, $\pi^{(j)}+\kappa$, $\pi^{(j)}+\iota+\kappa$
do not belong to ${\Pi} \cup \{\iota,\kappa\}$.
So, we have
$w_{\pi^{(j)}}+w_{\pi^{(j)}+\iota}+w_{\pi^{(j)}+\kappa}+w_{\pi^{(j)}+\iota+\kappa} = 1+0+0+0=1$,
and, by Lemma~\ref{l:RR}, $\bar w \not\in <R_\iota,R_\kappa>$.

By a similar argument, neither $R_\iota+ \zeroext{\iota} + \bar e^{(\iota)}$
nor $R_\kappa+ \zeroext{\kappa} + \bar e^{(\kappa)}$ contains $0^m$.
\proofend

\begin{theorem}
  Let $C\subset F^m$ be a $1$-code containing $0^m$; put $\dot C := C \setminus \{0^m\}$.
  Then the set
 $$ P(C) := \left( H \setminus \bigcup_{\iota\in \dot C}(R_\iota+ \zeroext{\iota} + \bar e^{(\iota)})
 \right) \cup \bigcup_{\iota\in \dot C}(R_\iota+ \zeroext{\iota})$$
 is a $1$-perfect code in $F^n$. Moreover,
\begin{equation}\label{eq:CtoP}
  C= \{ \iota \in F^m | (\iota,0^{n-m})\in P(C)\}.
\end{equation}
\end{theorem}
\proofr
We note that, by (\ref{eq:H}),  $\zeroext{\iota} + \bar e^{(\iota)}$
belongs to $H$ for all $\iota$; so, $R_\iota+ \zeroext{\iota} + \bar e^{(\iota)} \subset H$ for all $\iota$.

By Lemma~\ref{l:disj}, the sets $R_\iota+ \zeroext{\iota} + \bar e^{(\iota)}$, $\iota\in \dot C$,
are mutually disjoint.
Since they are subsets of a $1$-perfect code, their neighborhoods are also mutually
disjoint. Taking into account Lemma~\ref{l:O}, we see that $P(C)$ is a $1$-perfect code
by the definition.

It is easy to see that\\
(*) \emph{the only word in $H$ that has the form $(\alpha,0^{n-m})$
is the all-zeroes word.}

Furthermore, only $\zeroext{\iota}$ has such form in $R_\iota+ \zeroext{\iota}$, i.e.,\\
(**) \emph{if for some $\kappa\in F^m$
we have $(\kappa,0^{n-m}) \in R_\iota+ \zeroext{\iota}$,
then $\kappa = \iota$.}
Indeed, assume that $\zeroext{\kappa}=(\kappa,0^{n-m})\in R_\iota+ \zeroext{\iota}$.
Then $\zeroext{\kappa} + \zeroext{\iota} \in R_\iota \subset H$.
By (*), we have $\zeroext{\kappa} + \zeroext{\iota}=0^n$,
which proves the claim (**).

From (*) and (**) we conclude that (\ref{eq:CtoP}) is implied by the definition of $P(C)$.
\proofend

\begin{remark}
It is not difficult to see that ${\mathrm{rank}}(P(C))={\mathrm{rank}}(H)+{\mathrm{rank}}(C)$, where the rank of a code is the dimension
  of its linear span.
  So, if $m$ is sufficiently large, we can construct a code $P(C)$ of any rank from $2^m-m$ to $2^m-1$.
 This gives  (for large lengths) an interesting treatment of the result \cite[Proposition 6.2]{EV:94}
 on the existence of $1$-perfect
 codes of any ${\mathrm{rank}}\geq {\mathrm{rank}}(H)$.
\end{remark}
\begin{remark} As proved in \cite{Mal:2001:nonsyst}, it is enough to switch only
  seven disjoint cosets $R_\iota+z_\iota$,  $R_\kappa+z_\kappa$,  $R_\lambda+z_\lambda$,
  $R_{\iota+\kappa}+z_{\iota+\kappa}$, $R_{\iota+\lambda}+z_{\iota+\lambda}$,
  $R_{\kappa+\lambda}+z_{\kappa+\lambda}$, $R_{\iota+\kappa+\lambda}+z_{\iota+\kappa+\lambda}$
  in $H$ for some linearly independent $\iota$, $\kappa$, $\lambda$ and some
  $z_\iota,z_\kappa,z_\lambda,z_{\iota+\kappa},z_{\iota+\lambda},z_{\kappa+\lambda},z_{\iota+\kappa+\lambda}\in H$
  to obtain a nonsystematic $1$-perfect code. So, starting from $m=6$, we
  can construct a nonsystematic $1$-perfect code $P(C)$ with $|C|=8$.
\end{remark}

It would be interesting to define the minimal value of $n(m)$ such that any $1$-code of length $m$
can always be embedded in a $1$-perfect code of length $n(m)$. Recently, the similar question for
Steiner triple systems was solved \cite{BryHor:Lindner}:
any partial Steiner triple system of order $u$ can always be embedded in a Steiner triple systems
of order $v$ if and only if  $v\geq 2u+1$ and $v \equiv 1,3 \bmod 6$
(the last is the necessary and sufficient condition for the existence
of Steiner triple systems of order $v$). Taking into account the relation between
perfect codes and Steiner triple systems discussed in the introduction, we can state
that $n(m)\geq 2m+1$; the Theorem gives $n(m)\leq 2^m-1$.

The authors would like to thank the referees for their useful comments.

\providecommand\href[2]{#2} \providecommand\url[1]{\href{#1}{#1}}
  \def\DOI#1{{\small {DOI}:
  \href{http://dx.doi.org/#1}{#1}}}\def\DOIURL#1#2{{\small{DOI}:
  \href{http://dx.doi.org/#2}{#1}}}\providecommand\bbljun{June}

\end{document}